OpenStudy (anonymous):

A truck with a mass of 1.5 x 103 kg accelerates to a speed of 18.0 m/s in 12.0 s from a dead stop. Assume that the force of resistance is a constant 400.0 N during the acceleration. What is the average power developed by the truck’s engine?

OpenStudy (loser66):

@dan815

OpenStudy (dan815):

you need to find the total acceleration, if it took 12 seconds to get to 18.0 m/s velocity

OpenStudy (dan815):

err speed*

OpenStudy (dan815):

dist=1/2at^2 velocity=a*t 18/12=a 3/2= a a=3/2 m/s^2

OpenStudy (dan815):

for the truck to have a total acceleration of 3/2 m/s^2 we have to find the engines acceleration that is overcoming the deacceleration due to friction

OpenStudy (dan815):

acceleration due to friction = F= MA 400 = 1500 * A a=400/1500 <--- acceleration due to friction Therefore the total acceleration the engine must generate is 3/2 + 4/15 = (45+8)/30 =53/30 m/s^2

OpenStudy (dan815):

uhh so Power= Force * velocity

OpenStudy (dan815):

power=work/time

OpenStudy (dan815):

power=rate of work done

OpenStudy (dan815):

work = Force * distance so power= work / time power = Force * distance/time, but distance/time = speed = Force * speed

OpenStudy (dan815):

speed=3/2*t